Solve the system by substitution.
x + 4y = –4
3x – 3y = –12

Answer by checkley71(8403)

X+4Y=-4
X=-4Y-4 REPLACE X BY (-4Y-4) IN THE SECOND EQUATION
3(-4Y-4)-3Y=-12
-12Y-12-3Y=-12
-15Y=0
Y=0 SOLUTION
X=-4*0-4
X=-4 SOLUTION